# MGPATH: Vision-Language Model with Multi-Granular Prompt Learning for Few-Shot WSI Classification

### v3

@@ 4.2 Comparison to State-of-the-Art
, ABMIL Ilse et al. (2018), CLAM Lu et al. (2021), TransMIL Shao et al. (2021), DSMIL Li et al. (2021), GTMIL Zheng et al. (2022), DTMIL Zhang et al. (2022a), RRT-MIL Tang et al. (2024) and IBMIL Lin et al. (2023), and vision-language methods, including CoOp Zhou et al. (2022b), CoCoOp Zhou et al. (2022a), Metaprompt Zhao et al. (2024), TOP Qu et al. (2024a), ViLa-MIL Shi et al. (2024), MSCPT Han et al. (2024), QUILT Ikezogwo et al. (2024), CONCH Lu et al. (2024). Among these, QUILT and CONCH are foundation VLMs.
We provide different versions for our MGPath including CLIP backbone ReNet-50 (CLIP50) for TCGA-NSCLC and TCGA-RCC and ViT-16 backbone for TCGA-BRCA. We also provide a version using PLIP backbone, as well as GigaPath-PLIP, which was pre-trained on the Pathology dataset.
@@
As shown in Tables 2 and 1, our MGPath, based on CLIP50 and PLIP, outperforms several baseline models and achieves significant improvements over other VLMs with similar architectures, such as ViLa and MSCPT. The performance gain is particularly notable with the PLIP backbone. For example, on TCGA-BRCA using CLIP (ViT), MGPath achieves an accuracy of 77.10%, compared to 75.82% for MSCPT and 75.01% for ViLA-MIL. Additionally, with the PLIP backbone, MGPath surpasses MSCPT and ViLa-MIL by margins of approximately 3.5% to 5%.
GigaPath-PLIP is a strong pre-trained VLM.
We validated our whole-slide vision foundation model, pre-trained on 1.3 billion pathology images, using the PLIP text encoder. By incorporating pathology-specific features from Prov-GigaPath Xu et al. (2024), the integrated MGPath (PLIP-G) model demonstrated strong performance across multiple metrics on the TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA datasets. When compared to other foundation VLMs such as CONCH and QUILT, our model consistently outperforms them. For example, we achieve a 3% improvement in AUC over CONCH on both the TCGA-BRCA and TCGA-NSCLC datasets.
Table 2: Comparison of methods on TCGA-NSCLC, and TCGA-RCC datasets with few-shot settings. Results are shown for AUC, F1, and Accuracy (ACC).
@@
Max-pooling 197K 53.0$\pm$6.0 45.8$\pm$8.9 53.3$\pm$3.4 67.4$\pm$4.9 46.7$\pm$11.6 54.1$\pm$4.8
Mean-pooling 197K 67.4$\pm$7.2 61.1$\pm$5.5 61.9$\pm$5.5 83.3$\pm$6.0 60.9$\pm$8.5 62.3$\pm$7.4
ABMIL Ilse et al. (2018) 461K 60.5$\pm$15.9 56.8$\pm$11.8 61.2$\pm$6.1 83.6$\pm$3.1 64.4$\pm$4.2 65.7$\pm$4.7
CLAM-SB Lu et al. (2021) 660K 66.7$\pm$13.6 59.9$\pm$13.8 64.0$\pm$7.7 90.1$\pm$2.2 75.3$\pm$7.4 77.6$\pm$7.0
CLAM-MB Lu et al. (2021) 660K 68.8$\pm$12.5 60.3$\pm$11.1 63.0$\pm$9.3 90.9$\pm$4.1 76.2$\pm$4.4 78.6$\pm$4.9
TransMIL Shao et al. (2021) 2.54M 64.2$\pm$8.5 57.5$\pm$6.4 59.7$\pm$5.4 89.4$\pm$5.6 73.0$\pm$7.8 75.3$\pm$7.2
DSMIL Li et al. (2021) 462K 67.9$\pm$8.0 61.0$\pm$7.0 61.3$\pm$7.0 87.6$\pm$4.5 71.5$\pm$6.6 72.8$\pm$6.4
GTMIL Zheng et al. (2022) N/A 66.0$\pm$15.3 61.1$\pm$12.3 63.8$\pm$9.9 81.1$\pm$13.3 71.1$\pm$15.7 76.1$\pm$12.9
DTMIL Zhang et al. (2022a) 986.7K 67.5$\pm$10.3 57.3$\pm$11.3 66.6$\pm$7.5 90.0$\pm$4.6 74.4$\pm$5.3 76.8$\pm$5.2
IBMIL Lin et al. (2023) N/A 69.2$\pm$7.4 57.4$\pm$8.3 66.9$\pm$6.5 90.5$\pm$4.1 75.1$\pm$5.2 77.2$\pm$4.2
ViLa-MIL Shi et al. (2024) 8.8M/47M 74.7$\pm$3.5 67.0$\pm$4.9 67.7$\pm$4.4 92.6$\pm$3.0 78.3$\pm$6.9 80.3$\pm$6.2
CONCH (Lu et al. (2024)) 110M 89.46$\pm$10.2 78.5$\pm$9.31 78.78$\pm$9.1 88.08$\pm$4.59 78.21$\pm$4.2 71.67$\pm$19.4
QUILT Ikezogwo et al. (2024) 63M 79.66$\pm$13.19 72.30$\pm$13.35 72.42$\pm$13.24 96.92$\pm$1.6 78.46$\pm$5.55 86.34$\pm$1.56
MGPath (CLIP) 1.6M/39M 77.2$\pm$1.3 70.9$\pm$2.0 71.0$\pm$2.1 92.1 $\pm$ 2.8 76.5 $\pm$ 5.2 81.7 $\pm$ 2.9
@@
GigaPath-PLIP achieves competitive performance in zero-shot tasks.
@@ 4.4 Ablation Studies
We validate the use of Prov-GigaPath and PLIP under the following settings: (i) using full vision-language PLIP model; (ii) combining Prov-GigaPath with PLIP through the MLP layer which was pre-trained on the large-scale dataset; (iii) integrating Prov-GigaPath with PLIP through adaptor layers which were randomly initialized; (iv) utilizing Prov-GigaPath and an adaptor layer to map to the class output and only train MLP and last FFN layer of slide encoder; (v) only using Prov-GigaPath and an MLP layer to map to the class output and train MLP, the query matrix of last layer and the last FFN layer of slide encoder. Table 6 shows that using Prov-GigaPath combined with PLIP boosts the final performance compared to only using PLIP or Prov-GigaPath.
Table 4: Ablation studies on multi-granular (M-Gran), ratio combines two attention levels ($\alpha$), and message passing network types.
@@
MGPath (CLIP) 76.2$\pm$2.2 69.0$\pm$3.5 69.3$\pm$2.8
- w/o M-Gran (CLIP) 74.6$\pm$2.2 67.8$\pm$2.4 67.8$\pm$2.5
MGPath (PLIP-G) 91.7$\pm$3.6 84.2$\pm$4.6 84.4$\pm$4.5
- w/o M-Gran (PLIP-G) 90.6$\pm$4.5 82.4$\pm$5.7 82.5$\pm$5.7
MGPath, $\alpha=0.2$ 76.2$\pm$2.2 69.0$\pm$3.5 69.3$\pm$2.8
- $\alpha=0.5$ 73.7$\pm$3.1 67.4$\pm$2.6 67.8$\pm$2.7
- $\alpha=0.8$ 72.2$\pm$5.2 66.4$\pm$5.5 66.8$\pm$5.2
@@
MGPath (CLIP) 92.1$\pm$2.8 76.5$\pm$5.2 81.7$\pm$2.9
- w/o M-Gran (CLIP) 91.6$\pm$3.5 72.3$\pm$6.4 80.2$\pm$4.4
MGPath (PLIP-G) 98.1$\pm$0.6 85.7$\pm$1.1 89.9$\pm$2.0
- w/o M-Gran (PLIP-G) 98.1$\pm$0.6 85.0$\pm$4.0 89.3$\pm$3.0
@@
MGPath (PLIP) 592K 83.6$\pm$4.5 76.41$\pm$4.8 76.5$\pm$4.8
MGPath (PLIP-G) 5.35M 91.7$\pm$3.6 84.2$\pm$4.6 84.4$\pm$4.5
MGPath Random Adaptors 5.35M 91.4$\pm$4.2 82.8$\pm$5.7 83.0$\pm$5.6
GiGaPath Tuning (MLP + last FFN) 4.7M 62.7$\pm$3.5 64.66$\pm$5.3 52.8$\pm$3.4
GiGaPath Tuning (MLP + last Q-ViT) 5.8M 83.1$\pm$6.9 74.3$\pm$7.5 75.8$\pm$6.1
Table 6: Contribution of OT and multiple descriptive text prompts
@@
MGPath (OT, 4 text prompts) 76.2$\pm$2.2 69.0$\pm$3.5 69.3$\pm$2.8
MGPath (OT, 2 text prompts) 77.2$\pm$1.3 70.9$\pm$2.0 71.0$\pm$2.1
MGPath (Cosine, 2 text prompts) 75.8$\pm$3.7 68.3$\pm$4.5 68.4$\pm$4.5
@@
MGPath (OT, 4 text prompts) 92.1$\pm$2.8 76.5$\pm$5.2 81.7$\pm$2.9
MGPath (OT, 2 text prompts) 92.1$\pm$2.6 75.6$\pm$3.9 80.4$\pm$2.4
MGPath (Cosine, 4 text prompts) 91.8$\pm$2.8 75.9$\pm$4.3 80.5$\pm$2.6
Multi-Granular Prompt Learning.
In Table 6, we show the performance of MGPath with and without multi-granular (M-Gran) for CLIP (row 1 and 2) and PLIP-G (row 3 and 4) on TCGA-NSCLC dataset. It shows that using M-Gran improves the final performance of MGPath. This also happens on TCGA-RCC dataset. Table 6 also shows the impact of ratio when combining attention with graph and attention no graph on TCGA-NSCLC. It shows that with a ratio of 0.2/0.8 (0.2 for spatial attention obtained from graph structure and 0.8 for prototype-guided attention), MGPath achieves the highest performance.
OT as Alignment between Contextual Prompts.
Table 6 validates the use of OT in our MGPath on TCGA-NSCLC and TCGA-RCC. We see that using OT helps to boost the performance of MGPath (rows 1 and 2) compared to the use of cosine (rows 3 and 4). It also shows that the number of prompt vectors depends on each dataset. In the appendix, we also run with another version using unbalanced optimal transport (UoT). We observe that both UoT and OT provide good alignment quality, with UoT slightly outperforming OT. However, this advantage comes at the cost of increased running time.
@@ 4.5 Discussion
While we demonstrate significant improvements in few-shot and zero-shot WSI classification across several settings, this paper does not explore other important challenges. For example, how can we scale the current attention mechanism to handle even larger image patches (e.g., using Flash Attention Dao et al. (2022)), or extend the model from classification to tumor segmentation tasks Khened et al. (2021). Additionally, the potential for extending GiGaPath to integrate with other large-scale VLM models, such as CONCH Lu et al. (2024), remains unexplored.
@@ 5 Conclusion
High-resolution WSI is crucial for cancer diagnosis and treatment but presents challenges in data analysis. Recent VLM approaches, which utilize few-shot and weakly supervised learning, have shown promise in handling complex whole-slide pathology images with limited annotations. However, many overlook the hierarchical relationships between visual and textual embeddings, ignoring the connections between global and local pathological details or relying on non-pathology-specific pre-trained models like CLIP. Additionally, previous metrics lack precision in capturing fine-grained alignments between image-text pairs. To address these gaps, (i) we propose MGPath, which integrates Prov-GigaPath with PLIP, cross-aligning them with 923K domain-specific image-text pairs. (ii) Our multi-granular prompt learning approach captures hierarchical tissue details effectively, (iii) while OT-based visual-text distance ensures robustness against data augmentation perturbations. Extensive experiments on three cancer subtyping datasets demonstrate that MGPath achieves state-of-the-art results in WSI classification. We expect that this work will pave the way for combining large-scale domain-specific models with multi-granular prompt learning and optimal transport to enhance few-shot learning in pathology.
@@ 6 Acknowledgement
This work is supported in part by funds from the German Ministry of Education and Research (BMBF) under grant agreements No. 01D2208A and No. 01KD2414A (project FAIrPaCT).
@@
The authors gratefully acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy/Grete at NHR-Nord@Göttingen as part of the NHR infrastructure. The calculations for this research were conducted with computing resources under the project nim00014.
The authors gratefully acknowledge the computing time granted by the KISSKI project. The calculations for this research were conducted with computing resources under the project kisski-umg-fairpact-2.
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Duy M. H. Nguyen. Duy M. H. Nguyen and Daniel Sonntag are also supported by the XAINES project (BMBF, 01IW20005), No-IDLE project (BMBF, 01IW23002), and the Endowed Chair of Applied Artificial Intelligence, Oldenburg University.
Supplement to “MGPATH: Vision-Language Model with Multi-Granular Prompt Learning for Few-Shot WSI Classification”
@@ Appendix A Description of Dataset Splitting
TCGA-BRCA. This dataset contains 1056 whole slide images of breast invasive carcinoma. To conduct fair experiments, we adapted training and testing slides provided by the GitHub repository of MSCPT Han et al. (2024). In the MSCPT setup, 20% of the dataset was allocated for training, while the remaining 80% (833 slides) served as the test set. A fixed set of 16-shot WSIs was randomly sampled from the training set. Additionally, MSCPT specified the exact training and testing slides used in its experiments. However, there are 35 slides in which we got errors in the pre-processing steps; thus, we replaced those slides with the other ones (same number of WSI per class) downloaded from Cancer Genome Atlas (TCGA) Data Portal (GDC) The Cancer Genome Atlas () (TCGA).
TCGA-RCC &\&& TCGA-NSCLC. We adopt the same data splitting as in ViLa-MIL Shi et al. (2024), using 16-shot samples for training in each dataset. For testing, 192 samples were used for TCGA-RCC and 197 samples were used for TCGA-NSCLC.

### v5

@@ 4.2 Comparison to State-of-the-Art
, ABMIL (Ilse et al., 2018), CLAM (Lu et al., 2021), TransMIL (Shao et al., 2021), DSMIL (Li et al., 2021), GTMIL (Zheng et al., 2022), DTMIL (Zhang et al., 2022a), RRT-MIL (Tang et al., 2024) and IBMIL (Lin et al., 2023), and vision-language methods, including CoOp (Zhou et al., 2022b), CoCoOp (Zhou et al., 2022a), Metaprompt (Zhao et al., 2024), TOP (Qu et al., 2024a), ViLa-MIL (Shi et al., 2024), MSCPT (Han et al., 2024), QUILT (Ikezogwo et al., 2024), CONCH (Lu et al., 2024). Among these, QUILT and CONCH are foundation vision-language models (FVM).
Figure 4: Comparison of methods on TCGA-BRCA with 16-shot settings.
CLIP ImageNet Pretrained Methods # Param. TCGA-BRCA
AUC F1 ACC
[…]
4.3 MGPATH versions
We release four MGPATH variants. The ResNet-50 backbone CLIP-based models (MGPATH(CLIP)) for TCGA-NSCLC and TCGA-RCC, and a ViT-B/16 backbone (MGPATH(ViT)) for TCGA-BRCA. A pure PLIP version (MGPATH(PLIP)) is also provided. Finally, MGPATH(PLIP-G) configuration-combining the PLIP text encoder with the vision encoder from Prov-GigaPath pretrained on large-scale pathology dataset-is employed for all datasets.
@@ 4.3 MGPATH versions
As shown in Tables 1 and 4, our MGPATH(CLIP), MGPATH(ViT), and MGPATH(PLIP) outperforms several baseline models and achieves significant improvements over other VLMs with similar architectures, such as ViLa and MSCPT, under 16-shots setting. The performance gain is particularly notable with the PLIP backbone. For example, on TCGA-BRCA, MGPATH(ViT) achieves an accuracy of 77.10%, compared to 75.82% for MSCPT and 75.01% for ViLA-MIL. Additionally, with the PLIP backbone, MGPATH(PLIP) surpasses MSCPT and ViLa-MIL by margins of approximately 3.5% to 5%.
MGPATH(PLIP-G) is a strong pre-trained VLM.
By incorporating pathology-specific features from Prov-GigaPath (Xu et al., 2024), which is pre-trained on 1.3 billion pathology images, and PLIP text encoder, MGPATH (PLIP-G) model demonstrated strong performance across multiple metrics on the TCGA-NSCLC, TCGA-RCC, and TCGA-BRCA datasets. When compared to other foundation VLMs such as CONCH and QUILT, our model consistently outperforms them. For example, we achieve a 3% improvement in AUC over CONCH on both the TCGA-BRCA and TCGA-NSCLC datasets.
Table 1: Comparison of methods on TCGA-NSCLC, and TCGA-RCC datasets with 16-shots settings.
@@
[…]
@@
MGPATH(PLIP-G) achieves competitive performance in zero-shot tasks.
@@
Evaluating MGPATH(PLIP-G) with fewer Samples.
To assess the impact of number of training samples (number of shots $K$) on MGPATH(PLIP-G), we conducted experiments under 2-, 4-, 8-, 16-shots on TCGA-BRCA. Figure 5 presents box plots - each summaries five independent runs - that track F1 scores. Overall, MGPATH(PLIP-G) consistently surpassed the baselines across 2, 4, 8, and 16 shots.
Figure 5: F1 scores of MGPATH(PLIP-G) under 2-, 4-, 8-, and 16-shots training on TCGA-BRCA.
@@ 4.5 Ablation Studies
Multi-Granular Prompt Learning.
In Table 6, we show the performance of MGPATH(CLIP) and MGPATH(PLIP-G) with and without multi-granular on two datasets TCGA-NSCLC and TCGA-RCC. It shows that using M-Gran improves the final performance of MGPATH(CLIP) and MGPATH(PLIP-G) on both datasets. We further investigate the impact of ratio when combining prototype-guided attention with graph spatial attention on TCGA-NSCLC. Table 6 shows that with a ratio of 0.2/0.8 (0.2 for spatial attention obtained from graph structure and 0.8 for prototype-guided attention), MGPATH(CLIP) achieves the highest performance.
@@
We validate the use of Prov-GigaPath and PLIP under the following settings: (i) using full vision-language PLIP model; (ii) combining Prov-GigaPath with PLIP through the MLP layer which was pre-trained on the large-scale dataset; (iii) integrating Prov-GigaPath with PLIP through adaptor layers which were randomly initialized; (iv) utilizing Prov-GigaPath and an adaptor layer to map to the class output and only train MLP and last FFN layer of slide encoder; (v) only using Prov-GigaPath and an MLP layer to map to the class output and train MLP, the query matrix of last layer and the last FFN layer of slide encoder. Table 6 shows that using Prov-GigaPath combined with PLIP (MGPATH(PLIP-G)) boosts the final performance compared to only using PLIP (MGPATH(PLIP)) or Prov-GigaPath.
OT as Alignment between Contextual Prompts.
Table 6 validates the use of OT in our MGPATH(CLIP) on TCGA-NSCLC and TCGA-RCC. We see that using OT helps to boost the performance of MGPATH(CLIP) compared to the use of cosine. It also shows that the number of prompt vectors depends on each dataset. In the appendix, we also run with another version using unbalanced optimal transport (UoT). We observe that both UoT and OT provide good alignment quality, with UoT slightly outperforming OT. However, this advantage comes at the cost of increased running time.
Table 3: Ablation studies on multi-granular (M-Gran) on the proposed models’ performance.
@@
[…]
@@
[…]
Table 4: Ablation studies investigate the effect of the ratio combines spatial attention ($\alpha$) and prototype-guided attention in MGPATH(CLIP).
Configurations TCGA-NSCLC
AUC F1 ACC
[…]
@@
[…]
Table 6: Contribution of OT and multiple descriptive text prompts.
@@
[…]
@@
[…]
Effect of Large Language Models.
To further validate the impact of different LLMs on MGPATH(PLIP-G)’s performance, we conducted experiments with text prompts generated by various LLMs, including GPT-4o (OpenAI, 2024), Grok-3 (xAI, 2025), GPT-o3 (OpenAI, 2025), Gemini-2.5-pro (Team et al., 2023), Deepseek-R1 (DeepSeek-AI, 2025), and Mistral Medium 3  (Mistral AI, 2024). The results in 16-shot settings on TCGA-NSCLC are represented in Table 7. The experimental results demonstrated that MGPATH(PLIP-G) consistently exceeded the baseline across all text prompt sources. The model achieved the best performance with Mistral Medium 3  (Mistral AI, 2024), which produced the highest AUC (93.95 ± 2.71), F1 score (88.53 ± 4.29), and accuracy (88.63 ± 4.26). This experiment demonstrated the robustness of the performance of the proposed method in various LLMs and highlighted the critical role of precise visual descriptions on model performance enhancement.
Table 7: Ablation experiments of input text prompts from different LLM
TCGA-NSCLC
Methods AUC F1 ACC
GPT-4o (OpenAI, 2024) $93.02\pm 2.99$ $84.64\pm 4.75$ $84.77\pm 4.67$
Grok-3 (xAI, 2025) $93.07\pm 3.56$ $50\pm 4.53 85.58$$58\pm 4.51$
GPT-o3 (OpenAI, 2025) $93.87\pm 3.31$ $86.19\pm 5.28$ $86.29\pm 5.26$
Gemini-2.5-pro (Team et al., 2023) $93.17\pm 3.12$ $86.72\pm 3.50$ $86.80\pm 3.48$
Deepseek-R1 (DeepSeek-AI, 2025) $93.40\pm 2.80$ $86.87\pm 4.68$ $87.01\pm 4.63$
Mistral Medium 3 (Mistral AI, 2024) $\mathbf{93.95\pm 2.71}$ $\mathbf{88.53\pm 4.29}$ $\mathbf{88.63\pm 4.26}$
@@ 4.6 Discussion
While we demonstrate significant improvements in few-shot and zero-shot WSI classification across several settings, this paper does not explore other important challenges. For example, how can we scale the current attention mechanism to handle even larger image patches (e.g., using Flash Attention (Dao et al., 2022)), or extend the model from classification to tumor segmentation tasks (Khened et al., 2021). Additionally, the potential for extending GiGaPath to integrate with other large-scale VLM models, such as CONCH (Lu et al., 2024), remains unexplored.
@@ 5 Conclusion
High-resolution WSI is crucial for cancer diagnosis and treatment but presents challenges in data analysis. Recent VLM approaches, which utilize few-shot and weakly supervised learning, have shown promise in handling complex whole-slide pathology images with limited annotations. However, many overlook the hierarchical relationships between visual and textual embeddings, ignoring the connections between global and local pathological details or relying on non-pathology-specific pre-trained models like CLIP. Additionally, previous metrics lack precision in capturing fine-grained alignments between image-text pairs. To address these gaps, (i) we propose MGPATH(PLIP-G), which integrates Prov-GigaPath with PLIP, cross-aligning them with 923K domain-specific image-text pairs. (ii) Our multi-granular prompt learning approach captures hierarchical tissue details effectively, (iii) while OT-based visual-text distance ensures robustness against data augmentation perturbations. Extensive experiments on three cancer subtyping datasets demonstrate that MGPATH and its variants achieve state-of-the-art results in WSI classification. We expect that this work will pave the way for combining large-scale domain-specific models with multi-granular prompt learning and optimal transport to enhance few-shot learning in pathology.
6 Broader Impact Statement
MGPATH is developed to investigate the feasibility of applying vision-language models (VLMs) to pathology under few-shot learning settings, simulating scenarios with extremely limited training samples, such as rare diseases. However, the model is not intended for clinical use as a diagnostic tool or decision support system. It is not designed to replace professional medical judgment or diagnosis.
@@ 7 Acknowledgement
This work is supported in part by funds from the German Ministry of Education and Research (BMBF) under grant agreements No. 01D2208A and No. 01KD2414A (project FAIrPaCT). The authors gratefully acknowledge the computing time granted by the KISSKI project. The calculations for this research were conducted with computing resources under the project kisski-umg-fairpact-2. The authors also acknowledge the computing time granted by the Resource Allocation Board and provided on the supercomputer Emmy/Grete at NHR-Nord@Göttingen as part of the NHR infrastructure. The calculations for this research were conducted with computing resources under the project nim00014.
@@
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Duy M. H. Nguyen. Duy M. H. Nguyen and Daniel Sonntag are also supported by the XAINES project (BMBF, 01IW20005), No-IDLE project (BMBF, 01IW23002), and the Endowed Chair of Applied Artificial Intelligence, Oldenburg University. Anh-Tien Nguyen was a member of the Ph.D. program "Genome Science" – International Max Planck Research School.
Appendix
@@ Appendix A Description of Dataset Splitting
TCGA-BRCA. This dataset contains 1056 whole slide images of breast invasive carcinoma. To conduct fair experiments, we adapted training and testing slides provided by the GitHub repository of MSCPT (Han et al., 2024). In the MSCPT setup, 20% of the dataset was allocated for training, while the remaining 80% (833 slides) served as the test set. A fixed set of 16-shot WSIs was randomly sampled from the training set. Additionally, MSCPT specified the exact training and testing slides used in its experiments. However, there are 35 slides in which we got errors in the pre-processing steps; thus, we replaced those slides with the other ones (same number of WSI per class) downloaded from Cancer Genome Atlas (TCGA) Data Portal (GDC) (The Cancer Genome Atlas (), TCGA).
TCGA-RCC &\& TCGA-NSCLC. We adopt the same data splitting as in ViLa-MIL (Shi et al., 2024), using 16-shot samples for training in each dataset. For testing, 192 samples were used for TCGA-RCC and 197 samples were used for TCGA-NSCLC.